A line has a rise of 6 and a slope of 1/20. What is the run?

Answer by nerdybill(7384)

A line has a rise of 6 and a slope of 1/20. What is the run?
.
Since slope = rise/run
.
6/run = 1/20
cross-multiplying
6*20 = run
120 = run